OpenStudy (atomicreaper):

Find the remainder of (h^4 + h^2 – 2) ÷ (h + 3)

OpenStudy (amistre64):

P(x) = (x-c) Q(x) + R(x) for what value of x can we determine the value of R(x)?

OpenStudy (jdoe0001):

hint: (h+3) <= use synthetic division

OpenStudy (atomicreaper):

Uh what? I know to use synthetic division, thats the unit im studying. I hate division. See

OpenStudy (amistre64):

what you have is similar to: P(x)/(x-c) = Q(x) + R(x)/(x-c) this is simply how we turn an improper fraction into a mixed numeral: 12/5 = 2 + 2/5 right? if we multiply thru by (x-c) we get: P(x) = (x-c) Q(x) + R(x) the remainder will be known when (x-c) Q(x) = 0

OpenStudy (atomicreaper):

Continue the explanation please.

OpenStudy (amistre64):

let x=c P(c) = (c-c) Q(c) + R(c) P(c) = 0 + R(c) in other words, the remainder is the value of the top portion, when the bottom portion goes to 0

OpenStudy (amistre64):

if the top and bottom go to zero, then we would also know that they have a common factor ... but that is not your question at the moment

OpenStudy (atomicreaper):

No it is not

OpenStudy (amistre64):

what is the value of the top, when (h+3) = 0 ?

OpenStudy (atomicreaper):

Top?

OpenStudy (amistre64):

\[a\div b\implies\frac ab\]

OpenStudy (atomicreaper):

H?